# CNNs in the Air via Reconfigurable Intelligent Surfaces

### 07 Mar 2026 v2

@@ II System model
According to the definition of convolution, it is impossible to directly implement CNNs OTA without transformation. Unlike convolution, matrix multiplication can be inherently realized via OTA transmission. Therefore, we transform the convolutional operation into a multiplication operation by rearranging the matrices. To clearly illustrate this concept, Fig. 2 presents a toy example in which $\bf X$ and $\bf W$ represent a $3\times 3$ input image matrix and a $2\times 2$ convolutional kernel matrix, respectively. The top part of Fig. 2 shows the standard convolutional operation with a stride of $1$ and no padding, while the bottom part demonstrates the equivalent multiplication matrix operation. Specifically, matrices $\bf W$ and $\bf Y$ are unfolded into vectors $\bf\bar{W}$ and $\bf\bar{Y}$, respectively, and rearranged through a piece-wise block vectorization approach. To accommodate practical hardware constraints, the numbers of transmit and receive antennas are set to $4$ and $1$, respectively, matching the unfolded matrix size of $4\times 1$. Thus, after data rearrangement, the convolutional operation can be physically realized. Although Fig. 2 illustrates the case of a single kernel, multiple kernels can be extended similarly.
@@ V Numerical results
In Fig. 3, we study transmit power $P_{\rm max}$ versus classification accuracy for $L=1$, $K=3~{\rm dB}$, and $M=100$. The “Upper bound” scheme denotes the digital-domain Conv2d without OTA computation. It can be observed that the classification accuracy of all the schemes except the “Upper bound” increases with $P_{\rm max}$. This is expected, as higher $P_{\rm max}$ reduces the detrimental impact of noise at the receiver, thereby gradually approaching the “Upper bound” performance. Additionally, it is observed that the Conv2d-based schemes consistently outperform the ConvSD-based schemes. This is because the ConvSD layer is a simplified version of the Conv2d layer with weaker feature extraction capabilities. Furthermore, Conv2d MISO consistently outperforms Conv2d MIMO. This is because the numbers of adjustments for the precoder and RIS are $C_{\rm in}C_{\rm out}$ and $C_{\rm out}$, respectively, whereas they are $C_{\rm in}$ and $1$ for the Conv2d MIMO scheme, meaning that the DoFs available for emulation in the former scheme are much larger. It should be noted that this observation does not hold for the ConvSD scheme. The ConvSD MISO scheme outperforms the ConvSD MIMO scheme only when the transmit power is low, i.e., below $5~{\rm dB}$, but performs worse when the transmit power exceeds $5~{\rm dB}$. This is because the ConvSD MISO scheme adjusts the precoder $C_{\rm in}$ times without adjusting the combiner, whereas the ConvSD MIMO scheme adjusts both the precoder and the combiner once, thus striking different balances.

### 12 Mar 2026 v3

@@ II System model
According to the definition of convolution, it is impossible to directly implement CNNs OTA without transformation. Unlike convolution, matrix multiplication can be inherently realized via OTA transmission. Therefore, we transform the convolutional operation into a multiplication operation by rearranging the matrices. To clearly illustrate this concept, Fig. 2 presents a toy example in which $\bf X$ and $\bf W$ represent a $3\times 3$ input image matrix and a $2\times 2$ convolutional kernel matrix, respectively. The top part of Fig. 2 shows the standard convolutional operation with a stride of 11 and no padding, while the bottom part demonstrates the equivalent multiplication matrix operation. Specifically, matrices $\bf W$ and $\bf Y$ are unfolded into vectors $\bf\bar{W}$ and $\bf\bar{Y}$, respectively, and rearranged through a piece-wise block vectorization approach. To accommodate practical hardware constraints, the numbers of transmit and receive antennas are set to 44 and 11, respectively, matching the unfolded matrix size of $4\times 1$. Thus, after data rearrangement, the convolutional operation can be physically realized. Although Fig. 2 illustrates the case of a single kernel, multiple kernels can be extended similarly.
@@ V Numerical results
In Fig. 3, we study transmit power $P_{\rm max}$ versus classification accuracy for $L=1$, $K=3~{\rm dB}$, and $M=100$. The “Upper bound” scheme denotes the digital-domain Conv2d without OTA computation. It can be observed that the classification accuracy of all the schemes except the “Upper bound” increases with $P_{\rm max}$. This is expected, as higher $P_{\rm max}$ reduces the detrimental impact of noise at the receiver, thereby gradually approaching the “Upper bound” performance. Additionally, it is observed that the Conv2d-based schemes consistently outperform the ConvSD-based schemes. This is because the ConvSD layer is a simplified version of the Conv2d layer with weaker feature extraction capabilities. Furthermore, Conv2d MISO consistently outperforms Conv2d MIMO. This is because the numbers of adjustments for the precoder and RIS are $C_{\rm in}C_{\rm out}$ and $C_{\rm out}$, respectively, whereas they are $C_{\rm in}$ and 11 for the Conv2d MIMO scheme, meaning that the DoFs available for emulation in the former scheme are much larger. It should be noted that this observation does not hold for the ConvSD scheme. The ConvSD MISO scheme outperforms the ConvSD MIMO scheme only when the transmit power is low, i.e., below $5~{\rm dB}$, but performs worse when the transmit power exceeds $5~{\rm dB}$. This is because the ConvSD MISO scheme adjusts the precoder $C_{\rm in}$ times without adjusting the combiner, whereas the ConvSD MIMO scheme adjusts both the precoder and the combiner once, thus striking different balances.